Calculate the mass of candle wax(C25H52) that must burn to produce 4.00L of carbon dioxide at 21.0°C and 98.8kpa
1
Expert's answer
2020-06-11T10:27:44-0400

PV = nRT

n = PV / RT

T = 21.0 + 273.15 = 294.15 K

n = (98.8x1000x4.00x10-3) / (8.314 x 294.15)

n = 0.1616 mol

Each mole of CO2 is produced from 1/25 mol of candle wax (C25H52):

0.1616/25 = 0.00646 mol

M (C25H52) = 352.68

m = 0.00646 x 352.68 = 2.28 g